OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the points (-2, 6) and (-1, 4).

OpenStudy (anonymous):

use rise/run to find slope: (6-4)/(-2--1) = 2/(-1) =-2 now use point slope form (y-4)=(-2)(x--1) (y-4)=(-2)(x+1) plug and chug (y-4)=(-2x)-1 y=-2x+3 and that's your answer^

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

you're welcome! medal?

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -3 and 3x less than or eqaul to 12

OpenStudy (anonymous):

A number line with an open circle on -1, shading to the left, and a closed circle on 4, shading to the right.

OpenStudy (anonymous):

make a new q, get a new answer